OpenStudy (anonymous):

Nancy wants to invest her 50000.She invests part of it in corporate bonds paying 15% interest per year and the other part in a bank certificate that pays 7% interest per year.How much money should she invest at each rate to yield 6000 interest per year?

OpenStudy (anonymous):

07m+.15(50000-m)=6000 7m+15(50000-m)=600000 7m+750000-15m=600000 8m=150000 m=18750 50000-m=31250 Nancy needs to invest $18,750 at 7% and $31250 at 15%..

OpenStudy (anonymous):

Let x be the amount invested in the 15% account. Then 50,000 - x is the amount invested in the 7% account. Total interest from both accounts is .15x + .07(50,000 - x) = 6,000 .08x + 3,500 = 6,000 .08x =2,500 x = 31,250 50,000 - x = 18,750. Nancy should invest %31,250 in the 15% account and $18,750 in the 7% account.

OpenStudy (anonymous):

thank you guys.

OpenStudy (anonymous):

:^)